5. A machine can be purchased with a $7,000 down payment, with 72 months of payments of $960.00. If the interest rate is 24% compounded monthly, what is the price of the machine?
Answer #1

Monthly (nominal) interest rate = 24%/12 = 2%

Loan amount ($) = Monthly payment x P/A(2%, 72) = 960 x 37.9841** = 35,504.74

Price of machine ($) = Loan amount + Down payment = 35,504.74 + 7,000 = 42,504.74

**From P/A Factor table